### v2

# VulDetectBench: Evaluating the Deep Capability of Vulnerability Detection with Large Language Models

Abstract
Large Language Models (LLMs) have training corpora containing large amounts of program code, greatly improving the model’s code comprehension and generation capabilities. However, sound comprehensive research on detecting program vulnerabilities, a more specific task related to code, and evaluating the performance of LLMs in this more specialized scenario is still lacking. To address common challenges in vulnerability analysis, our study introduces a new benchmark, VulDetectBench, specifically designed to assess the vulnerability detection capabilities of LLMs. The benchmark comprehensively evaluates LLM’s ability to identify, classify, and locate vulnerabilities through five tasks of increasing difficulty. We evaluate the performance of 17 models (both open- and closed-source) and find that while existing models can achieve over 80% accuracy on tasks related to vulnerability identification and classification, they still fall short on specific, more detailed vulnerability analysis tasks, with less than 30% accuracy, making it difficult to provide valuable auxiliary information for professional vulnerability mining. Our benchmark effectively evaluates the capabilities of various LLMs at different levels in the specific task of vulnerability detection, providing a foundation for future research and improvements in this critical area of code security. VulDetectBench is publicly available at https://github.com/Sweetaroo/VulDetectBench.
1 Introduction
Recent advancements in Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains. The training datasets of some of these LLMs encompass vast quantities of program code, or the models have been fine-tuned on specific code datasets, endowing LLMs with significant abilities[1] in code understanding, generation, and summarizing. This development leads to a deeper investigation into whether current LLMs possess the capacity for vulnerability detection within code and the extent of effectiveness they can achieve in this realm.
Vulnerability analysis is a complex and systematic task, with numerous traditional methods and methods using deep learning have been proposed. Traditional methods focus on static program analysis[2, 3]and dynamic testing[4, 5]to identify and locate vulnerabilities in programs, improving the accuracy of analysis results while reducing false positives, and improving analysis efficiency and scalability for large programs have been the focus of research in traditional methods. Previous work based on deep learning approaches [6] creates a dataset comprising programs and their associated vulnerabilities. It applies deep learning techniques to determine the presence of vulnerabilities in a program or predict their specific type. However, due to limitations in dataset size and quality, these methods struggle to handle cases that fall beyond the training data distribution, thus hindering effective application in real-world scenarios [7].
Currently, there has been research into the capabilities of LLMs in identifying and repairing vulnerabilities in code, and the creation of datasets to analyze the ability of LLMs to detect vulnerabilities (i.e. determine whether code has vulnerabilities). However, a benchmark for evaluating the vulnerability detection capabilities of LLMs and a comparative analysis of various known open source and proprietary LLMs against such a benchmark is lacking. Different LLMs, due to their different sizes in terms of parameters, have different capabilities in different subtasks of vulnerability detection. Using a reasonable benchmark to evaluate the performance of different LLMs in these subtasks can provide critical reference information for subsequent research.
The task of vulnerability detection presents new challenges for LLMs. Firstly, programs with vulnerabilities in real-world scenarios tend to be large, and LLMs process these programs purely as text. Directly inputting such extensive programs into LLMs and expecting them to identify vulnerabilities is a daunting task. Secondly, even if a model can determine the presence of vulnerabilities within a program, we also need to know the types of vulnerabilities present. Different types of vulnerabilities have varying conditions for being triggered or exploited, which can provide security researchers with critical information for prioritizing patching efforts. Moreover, it’s essential to ascertain whether the model can accurately locate the vulnerabilities, including the root causes and the specific locations where they are triggered. This capability significantly enhances the accuracy of vulnerability identification. These tasks become progressively more challenging, but the latter ones offer more valuable information for vulnerability analysis.
In this work, we construct a multi-task benchmark test centered around the specific concrete requirements of vulnerability analysis to evaluate the capability of the LLMs in vulnerability detection at multiple levels. We evaluate a total of 17 LLMs, including three closed-source models. The smallest model parameter with known parameters is 6B and the largest is 70B. As show in Figure 1, we designed 5 tasks of progressive difficulty to progressively analyze the performance of the LLMs in existence detection, type classification, and vulnerability location. Our benchmark includes both artificially constructed datasets and real-world datasets. Furthermore, the programs in our benchmark dataset are written in C/C++, as a vast number of open-source libraries and kernels are developed in C/C++, and due to issues such as memory management, high-risk vulnerabilities like memory leaks frequently occur in C/C++ programs. Therefore, the capability of LLMs to detect vulnerabilities in large-scale C/C++ programs is highly representative.
Our main contributions are threefold, 1): High-quality benchmark: We present a comprehensive benchmark meticulously crafted to evaluate the vulnerability detection abilities of large-scale models. The benchmark comprises a refined vulnerability dataset and incorporates five distinct assessment tasks, each focusing on different aspects of vulnerability analysis. 2): Comprehensive evaluations: We conducted a comprehensive evaluation of the vulnerability detection capabilities of 17 existing large language models across 5 tasks. 3)Thorough analysis and new findings: We performed an in-depth analysis of the evaluation results from various perspectives, shedding light on the strengths and limitations of existing large language models in vulnerability detection. Our findings lay the foundation for advancing the understanding and application of large language models in the domain of vulnerability detection.
2 Related Work
2.1 Software Vulnerability and Detection Method.
Vulnerability is a flaw or weakness in a software program that can be exploited by an attacker to perform unauthorized actions within a computer system. Vulnerability detection involves identifying weaknesses in software programs that can be exploited by threats to compromise the security, functionality, or management of the software system. Researchers have proposed a large number of methods for identifying vulnerabilities in programs, which can be divided into two main categories: static analysis vulnerability detection[2] and dynamic analysis methods[8].
Static analysis[9][10] typically relies on feature collection and templates to provide fast processing speeds. Different static analysis tools[11][12][13] can deal with different types of vulnerabilities. However, due to the lack of contextual information from the execution environment, it often results in a high rate of false positives[9], meaning that many vulnerability alerts are not true vulnerabilities, requiring manual verification of these alerts.
Dynamic analysis, such as fuzzing tests[14, 5], explores as many parts of a program as possible by constructing a large diversity of program inputs, while dynamic fuzzing tests require the execution of a program to accurately identify vulnerabilities by directly observing the behavior of the vulnerabilities during execution. However, dynamic analysis is difficult and computationally expensive[4] to explore all states of a large-scale program due to the complexity of the program, such as a large number of loops and branches. For security analysts, the operational costs are prohibitive.
2.2 Deep Learning in Vulnerability Detection
Deep learning methods[6] transform program code into code slices[15] or graphical structures[3], such as CFGs, DFGs, and PDGs, and train them using LSTM[6] or graph neural networks[3]. These methods offer higher accuracy based on the learned features, significantly reducing the complexity involved in traditional dynamic and static analyses. However, these methods are often limited by the effect of the models themselves and the datasets used. Performance can significantly decline when the distribution of vulnerabilities in the datasets do not match real-world scenarios. Studies[7] have shown that deep learning’s performance in practical settings can dramatically decrease due to issues with the datasets.
Large language models are trained on extensive datasets that include program codes, vulnerability descriptions, and numerous code commits from open-source libraries, providing a certain level of understanding of various forms of vulnerabilities. These models treat program code as textual input, thereby rapidly assisting in vulnerability analysis. Despite these capabilities, research[16] indicates that current large language models face limitations when dealing with complex programs in real-world scenarios. These models may not fully comprehend the complex logic and structure of advanced programs, necessitating further optimization and adjustment to adapt to real-world applications.
Recent studies also utilize LLMs for vulnerability detection, introducing new datasets[17][18] or benchmark[19] focused on binary classification[17][16] or multi-classification [19] of input program. However, these studies typically use function-level inputs, which do not accurately represent real-world vulnerability analysis scenarios, where vulnerabilities and their root causes, along with trigger points, often span multiple functions. Additionally, the sources of data used to construct benchmarks are crucial. Research indicates[18][20] that the quality of existing datasets varies, influenced by the collection methods and their authenticity, whether synthetically generated or derived from real-world scenarios. The quality of labeling also varies across these datasets. These factors must be considered to ensure the reliability and validity of benchmarks.
3 Benchmark Construction
Figure 2 shows the overall process of VulDetectBench from construction to evaluation, and next we describe the dataset construction and specific task design respectively.
Data Source
In this part, our goal is to construct a benchmark from high-quality datasets of code vulnerabilities. As shown in Table 1. We collect real-world datasets from nine popular open source projects and a synthetic dataset from the National Institute of Standards and Technology Software Assurance Reference Dataset (NIST SARD), which provides rich annotated data on code vulnerabilities. Each entry in the dataset includes a detailed classification, description and key information about the code associated with the formation and triggering of the vulnerability, facilitating the design of various tasks. We also include two open source datasets: Big-Vul[21] and Devign[22]. More detailed information about dataset can be found in Appendix.
| Dataset Name | Version | Project | Type | Entries | Accuracy | ||
| Apache Subversion [23] | 1.8.3 | Subversion | Real | 638 | 1.000 | ||
| Wireshark [24] | 1.10.2 | WireShark | Real | 637 | 1.000 | ||
| Tree [25] | 1.7.0 | Tree | Real | 380 | 1.000 | ||
| PostgreSQL [26] | 9.2.4 | PostgreSQL | Real | 637 | 1.000 | ||
| OpenSSL[27] | 1.0.1e | OpenSSL | Real | 636 | 1.000 | ||
| GNU Grep[28] | 2.14 | Grep | Real | 380 | 1.000 | ||
| FFmpeg [29] | 1.2.2 | FFmpeg | Real | 637 | 1.000 | ||
| Gimp[30] | 2.8.8 | GIMP | Real | 637 | 1.000 | ||
| Juliet C/C++ [31] |
SARD | Synthetic | 64099 | 1.000 | ||
| Devign [22] | - | FFmpeg, Qemu, Wireshark | Real | 27318 | 0.800 | ||
| Big-Vul [21] | - | Multiple sources | Real | 43523 | 0.543 |
Task Design
VulDetectBench is a comprehensive benchmark tailored to large language models for detecting vulnerabilities in code. The benchmark consists of five tasks of increasing difficulty that provide an in-depth assessment of the vulnerability detection capabilities of LLMs: Task 1 involves a binary classification to determine the presence of code vulnerabilities. Task 2 is a multi-class classification task aimed at identifying the CWE (Common Weakness Enumeration) classification of code vulnerabilities. Starting from Task 3, we increased the difficulty level to explore whether models can truly understand the specific content of vulnerabilities. Tasks 4 and 5 further assess the model’s ability to identify the root causes and trigger points of vulnerabilities, respectively. These tasks are critical for in-depth vulnerability analysis and represent the most challenging aspects, testing the models’ semantic understanding of program vulnerabilities comprehensively.
| No. | Task |
| 1 | Vulnerability Existence Detection | 1000 | 48 | 78 | 90 | 23 | 36 | 114 | 50 - 3493 | ||||||||
| 2 | Vulnerability CWE Type Inference | 500 | 48 | Same Above | 265 - 3372 | ||||||||
| 3 | Key Objects and Functions Identification | 100 | 38 | 476 | 191 | 88 | 78 | 89 | 1017 - 3269 | ||||||||
| 4 | Vulnerability Root Cause Location | 100 | 38 | Same Above | 1010 - 3262 | ||||||||
| 5 | Vulnerability Trigger Point Location | 100 | 38 | Same Above | 1011 - 3363 |
As shown in Table 2. To enhance the precision of assessing the vulnerability analysis capabilities of Large Language Models (LLMs), the benchmarks for Tasks 1 and 2 are designed to contain identical datasets, each comprising 1,000 data entries. Similarly, Tasks 3, 4, and 5 are standardized with exactly the same sample data, encompassing 100 data entries for each task. This uniformity allows for direct comparison across tasks and facilitates tracing specific samples to determine the maximum difficulty level the models can handle.
Task 1: Vulnerability Existence Detection
For Task 1, we examine the capability of existing LLMs to perform the relatively simple binary classification task of determining the presence of vulnerabilities in a program. Vulnerable samples for this task are constructed from both real-world and synthetically generated datasets that are well-annotated and provide comprehensive relevant information. The test cases in task1 require the LLMs to output a "YES" or "NO" to indicate the existence of a vulnerability. During evaluation, the accuracy and F1 score of the models are assessed based on whether their output matches the correct answer.
Task 2: CWE Type Inference
Specifically, for this task, we present five different options for each sample: the optimal answer, which is the sample’s actual CWE type, and a suboptimal answer derived from the CWE-1000 VIEW hierarchy[32], representing an ancestor node within four layers. Additionally, three incorrect choices are provided: two unrelated CWE types and one "No Vulnerability" option. Models must correctly identify the CWE type or select the correct option letter, with scoring based on two systems. The Moderate Evaluation (ME) awards 1 point for the optimal choice and 1 point for choosing the suboptimal without the optimal, whereas the Strict Evaluation (SE) grants 1 point for the optimal and 0.5 points for the suboptimal, with no points for other selections. This method scrutinizes the models’ depth of understanding and accuracy in inference vulnerability types within complex C/C++ program structures.
Task 3: Key Data Objects and Functions Identification
Identifying key data objects and function calls associated with vulnerabilities is crucial in vulnerability analysis. For instance, analyzing CWE 476 buffer overflow vulnerabilities involves examining the buffer and related memory read-write functions. LLMs’ ability to identify these key elements is essential for accurate vulnerability analysis. In Task 3, both Macro Recall (MAR) and Micro Recall (MIR) metrics are used. MIR is particularly effective in mitigating fluctuations caused by sparse labels in the ground truth, providing a more stable evaluation of model performance.
Task 4: Root Cause Location
For Task 4 in our benchmark, we focus on evaluating LLMs’ ability to locate the root cause of a vulnerability within a program’s codebase, a crucial aspect for precise vulnerability analysis. We use a data set where each vulnerability’s root cause is uniquely labeled, and employ a prompt designed to compel the model to identify the specific code region associated with the root cause. By extracting these identified regions, we calculate the recall of the model’s responses, thus measuring how effectively it can pinpoint the root causes within the textual code data.
Task 5 : Trigger Point Location
In Task 5, we examine the LLMs’ capacity to identify the precise trigger point of a vulnerability, typically localized to a few lines or a single line of code. This task challenges the models to comprehend extensive code text and accurately determine the specific trigger point, which is essential for effective vulnerability localization. The task involves a similar setup to Task 4, with a prompt that requires the model to specify the code region during its response. The performance is evaluated by calculating the recall of the model’s output in identifying the exact lines or line where the vulnerability is triggered. For both Task 4 and Task 5, the primary metric for evaluation is line-of-code recall. This involves segmenting identified code snippets from the model’s output, analyzing them by lines, and assessing the recall of the lines that correspond to the standard answers, using a macro-mean recall for the final metric. For Task 5, trigger point code lines are extracted and analyzed similarly. Additionally, an alternative recall metric is computed—defined by the intersection over union of the model’s output and the true labels—to ensure a balanced assessment of the model’s accuracy and precision in both identifying and localizing vulnerabilities.
URS stands for Union Recall Score and IL stands for the intersection lines of code between LLMs’ output and the ground truth answers. ROL stands for Result Output Lines and UL stands for Union Lines between LLMs output and ground truth answers. ORS is designed to detect whether the model complies well with the instructions and only outputs codes, and on the other hand, to mitigate the large number of false alarms introduced by the model when it outputs a large number of codes to increase its score, which is difficult to judge.
4 Experiment
4.1 Setup
We selected 17 models from 10 different families in the experiment, including proprietary models like GPT-4, Google Gemini-Pro, and Ernie 4.0. The models varied in size, with parameters ranging from 6B to 70B. Mixtral 8 * 22B, an open-source model, had the highest parameter count. We ensured that at least two models with different parameter sizes are chosen from each family for comparison. Notably, Deepseek and CodeLlama underwent code-specific pre-training.
The computational resources are standardized using NVIDIA GeForce RTX 4090 graphics cards. Single RTX 4090 units are used for models up to 7B, while two RTX 4090s are used in parallel for larger models (13B-14B) to maintain optimal conditions and avoid performance bottlenecks.
4.2 Overall Results
As presented in Table 3, we evaluate the performance of 17 Large Language Models (LLMs) across five tasks, using two distinct metrics for each task. For Task 1, we measure Accuracy and F1 Score. In this task, Ernie 4.0, a proprietary model, outperforms all others, achieving the highest Accuracy of 85.01% and F1 Score of 86.65%. Among the open-source models, Mixtral-8*22B shows the best performance in Accuracy. For models with a smaller number of parameters, Meta-Llama-8B-instrcut demonstrates superior performance, achieving an F1 Score of 76.53%.
| Model | Size |
| GPT4 | - | 71.43 | 79.30 | 92.96 | 95.17 | 20.21 | 16.07 | 24.26 | 27.07 | 13.00 | 17.85 | ||||||||||
| ERNIE4.0 | - | 85.01 | 86.65 | 72.50 | 70.00 | 27.87 | 22.54 | 11.77 | 27.99 | 22.43 | 10.38 | ||||||||||
| Gemini-pro | - | 73.10 | 75.74 | 70.10 | 77.00 | 13.03 | 10.55 | 14.64 | 23.51 | 07.56 | 18.89 | ||||||||||
| Deepseek | 7B | 71.30 | 61.16 | 37.60 | 40.20 | 17.81 | 13.67 | 05.36 | 09.22 | 04.30 | 08.83 | ||||||||||
| Qwen | 7B | 62.30 | 43.31 | 60.10 | 62.20 | 13.63 | 11.15 | 04.34 | 10.09 | 05.56 | 10.95 | ||||||||||
| Qwen | 14B | 69.10 | 55.67 | 77.90 | 82.80 | 16.49 | 13.67 | 06.40 | 08.32 | 03.14 | 04.81 | ||||||||||
| ChatGLM3 | 6B | 69.90 | 65.37 | 39.90 | 46.60 | 00.16 | 00.12 | 00.19 | 01.33 | 00.42 | 01.12 | ||||||||||
| Vicuna | 13B | 57.60 | 31.17 | 67.30 | 74.40 | 07.30 | 06.35 | 0 | 0 | 0 | 0 | ||||||||||
| Vicuna | 7B | 48.60 | 65.27 | 31.30 | 42.00 | 07.79 | 06.24 | 0 | 0 | 0 | 0 | ||||||||||
| CodeLlama | 13B | 47.90 | 58.81 | 32.80 | 44.60 | 10.34 | 08.51 | 03.30 | 03.47 | 01.29 | 01.89 | ||||||||||
| CodeLlama | 7B | 36.40 | 53.37 | 36.70 | 41.60 | 06.28 | 05.04 | 01.69 | 02.55 | 00.69 | 01.31 | ||||||||||
| Llama3 | 8B | 69.40 | 76.53 | 62.30 | 68.40 | 22.83 | 17.99 | 00.19 | 01.10 | 00.17 | 00.53 | ||||||||||
| Llama2 | 7B | 47.90 | 64.19 | 41.40 | 54.40 | 11.47 | 09.59 | 0 | 0 | 0 | 0 | ||||||||||
| Llama2 | 13B | 70.37 | 73.67 | 58.70 | 67.40 | 10.99 | 09.23 | 00.50 | 01.20 | 00.10 | 00.20 | ||||||||||
| Llama3 | 70B | 47.45 | 60.33 | 26.00 | 17.00 | 09.18 | 07.43 | 0 | 0 | 0 | 0 | ||||||||||
| Mixtral | 8*7B | 76.42 | 79.00 | 62.00 | 58.40 | 21.61 | 17.51 | 01.66 | 04.28 | 05.74 | 03.51 | ||||||||||
| Mixtral | 8 *22B | 81.82 | 84.47 | 77.80 | 74.80 | 30.26 | 24.10 | 17.49 | 11.46 | 06.17 | 02.83 |
In Task 2, the performance of different LLMs varies significantly. Some models, such as GPT-4, Qwen-14B-Chat, and Vicuna-13B-v1.5, perform better than in Task 1, while the majority of the models exhibit a substantial decline in performance. It is important to note that, generally, using the Moderate metric leads to an improvement of approximately 5%, and in some cases, such as with Vicuna-7B-v1.5, the improvement reaches 12%. This suggests that the models’ ability to predict vulnerability types is influenced by their training data, resulting in varied performance under different precision requirements of the same task.
In Task 3, the open-source model Mixtral-8*22B demonstrates a significant advantage, achieving a Macro Recall of 30.26%. This performance level provides valuable insights for vulnerability analysis. Among the smaller models, Meta-Llama-8B-instruct also performs well, outperforming other open-source models by an average of more than 5% in Macro Recall.
For Tasks 4 and 5, we employ two different metrics for evaluation, as outlined in Equations 2. We observe a significant performance decline across all models in these tasks, yet the disparity in effectiveness between models remains substantial. Some open-source models, such as the Vicuna, Llama families, and ChatGLM3-6B, struggle considerably with these tasks, nearly failing to accomplish them. GPT-4 and Ernie 4.0 exhibit the best performance in both tasks. However, we note that Ernie 4.0 experiences a decline of over 10% in LOC Recall based on intersections, and upon analyzing the actual output, we find that Ernie 4.0’s responses contain a substantial amount of irrelevant content, which adversely affects practical vulnerability analysis. Based on this comprehensive analysis, GPT-4 emerges as the most effective model in these tasks.
Overall, we observe that existing LLMs excel in the binary classification task of detecting vulnerabilities in programs (Task 1), achieving best accuracy above 85% for inputs up to 4K tokens. This performance is consistent across both open-source and proprietary commercial models. However, their performance declines in Task 2, which involves identifying specific CWE vulnerability types. Furthermore, all models significantly underperform in tasks that require identifying key data objects and function calls associated with vulnerabilities, as well as in accurately pinpointing the root causes and trigger points of these vulnerabilities.
4.3 In-depth Analysis
Model Size Influence
In this section, we conduct a detailed analysis of models from the same family but of varying sizes on their performance across the aforementioned five tasks. Generally, it is observed that models with a larger number of parameters tend to perform better on these tasks. Specifically, we exemplify this trend using the models Mistral, Llama2, Qwen, and CodeLlama. As shown in Figure 3, for five LLM families, we analyze the impact of model parameter size on performance across different tasks. It is evident that models with larger parameters generally perform better across all tasks. In simpler tasks, the increase in parameter size correlates with greater performance improvements, with a maximum increase of 46.9% in Task 1 and up to 115.0% in Task 2. However, the extent of performance improvement varies under different metrics. For Task 1, the metric based on the F1 score is more stringent than the Accuracy metric, resulting in a lower percentage of improvement for the models, especially for the Vicuna-13b-v1.5, where the F1 Score decreased by 52.2% compared to the same family’s 7B model.
Semantic Understanding of Vulnerability
To better understand the extent of the LLMs’ understanding of vulnerabilities, we use 100 identical test cases from tasks one to five for analysis. Despite their ability to accurately determine the presence of vulnerabilities and even identify CWE types, this does not mean that the models truly understand the specific details of the vulnerabilities. For example, they are unable to accurately identify data objects and critical function calls that are closely associated with vulnerabilities, nor can they accurately locate root causes and trigger points. This suggests that large models assess the presence of vulnerabilities primarily from the perspective of high-level language features, and lack a deeper understanding of the specific mechanisms of vulnerability occurrence.
As shown in Figure 4.3. Through a horizontal comparison among different tasks assigned to models, it is evident that performance on simpler tasks does not fully reflect a model’s capabilities in detecting code vulnerabilities. Taking GPT4 as an example, its performance in Task 1 suggests it can effectively identify the presence of vulnerabilities in code. However, its reduced performance in Task 2 indicates a biased judgment in detecting code vulnerabilities during Task 1. Further analysis of its performance in Tasks 3, 4, and 5 reveals that it struggles to accurately locate the trigger points, root causes, and critical data types associated with code vulnerabilities. Therefore, we conclude that the current capabilities of LLMs in vulnerability detection are limited and should be evaluated based on tasks of varying difficulty to accurately assess their ability to detect code vulnerabilities.
Comparison with Traditional Tools
Vulnerability analysis of programs using LLM has the intuitive advantage of being less dependent on the integrity of the program. Traditional static program analysis tools such as Infer require compilation of the program, which transforms program syntax trees (ASTs), program control flow graphs (CFGs), and other forms of program representation. CodeQL[11] requires the use of manually constructed databases to perform vulnerability matching. This does not allow for rapid vulnerability analysis, and we choose a commonly used tool flawfinder[12] that can directly analyze portions of a program’s code for comparative experiments. This tool can complete five tasks in VulDetectBench, but none of them give correct results and produce a large number of false positives. More detailed descriptions are provided in the Appendix.
5 Conclusion, Limitations and Societal impacts
Conclusions We have developed a new benchmark for evaluating LLMs on different tasks, using high quality real data and long program files as input. Our design of tasks of varying difficulty shows that current LLMs, both open source and proprietary, perform well on basic vulnerability detection tasks and can serve as auxiliary tools for vulnerability analysis. For example, these models are effective at detecting the presence of vulnerabilities and inferring the CWE types of the vulnerabilities. However, for more specialized tasks such as identifying key data objects and determining the root cause of vulnerabilities, their performance is still insufficient to replace traditional tools or provide effective analytical support.
Limitations While our benchmark provides a comprehensive evaluation of LLMs on vulnerability detection tasks, there are opportunities for future work to expand the dataset to include more programming languages and a wider range of vulnerability types.
Societal impacts Our benchmark can facilitate the development of large language models for vulnerability detection. Meanwhile, Using large language models for vulnerability detection may lead to malicious exploitation and over-reliance, necessitating responsible development and oversight.